Word problems, please help.
1. During the NBA’s 2001-2002 regular season Sacramento won 2 games less than three times of the games they lost. This team played a total of 82 games in the regular season. How many games did Sacramento loose and win in their 2001-2002 regular season?
2. During the United States Presidential campaign of 2000, candidates Al Gore and George W. Bush accumulated a total of 537 electoral votes between the both of them. George W. Bush obtained 5 more votes than Al Gore. What percentages of votes did each of the presidential candidates obtain?
3. Francisco got a score of 90 and 82 points in the first two course evaluations respectively. The third assessment is counted as two scores. If Francisco wants to ensure a final grade of 90%, what is the minimum score he must obtain in his third assessment to ensure that final 90% score?

Answer by Boreal(15235)

lose=x
win=3x-2
they add to 82, so 3x-2+x=82; 4x=84, x=21 games lost, 61 games won. 21 is ANSWER.
------------------
B=x+5; G=x
2x+5=537
2x=532
x=266 G 49.53%
x+5=276 B 50.47%
-----------------
Four tests essentially, and he has taken 2. For an average of 90, the total score for the 4 must be 360.
right now he has a score of 172
That means he needs 188 for the final test, and divided into equal parts, that is a 94 overall.